OpenStudy (anonymous):

Use properties of logarithms to find the exact value of the expression 2logs5+log8-log2.

jhonyy9 (jhonyy9):

log5(5) =1 and log(2)=1 log8 = log2^3 =3 after this hope that you can do it

OpenStudy (anonymous):

yea my key says that the answer is 2?

jhonyy9 (jhonyy9):

why 2

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

ummm i have no idea..... i double checked the key but it says the answer was 2. a.-2 b.-1 c. 0 d. 1 e. 3 f. 2 it says the answer if F

jhonyy9 (jhonyy9):

the first term is 2 ,the second is 3 and the 4th is 1 so hence you get 2+3-1 =4

jhonyy9 (jhonyy9):

if the right 2 - so this you can get it if you log-ing the 4 because log(4) = 2log2 = 2

OpenStudy (anonymous):

hmm...so your saying because of log of 4...its 2?